OpenStudy (anonymous):

how to solve integral of 5xsinxcosx

OpenStudy (anonymous):

convert sinx cos x = 1/2 sin 2x first then put it back to integral let u = 2x and ..... done

OpenStudy (anonymous):

the leftover is not hard for you, I think so. of course it's not done as I state, but just some simple steps you can get the answer

OpenStudy (callisto):

I guess you'll need integration by parts.

OpenStudy (anonymous):

wat answer did u guys get

OpenStudy (callisto):

What have you got?

OpenStudy (anonymous):

massive answer

OpenStudy (callisto):

Type/write it here.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so

OpenStudy (anonymous):

@ramstar I don't have the answer, just wait for you.

OpenStudy (anonymous):

so let =5x/2 v =sin(2x)

OpenStudy (anonymous):

u'=5/2 and v' =2 cos (2x)

OpenStudy (anonymous):

then i used this formula uv-integral u'v

OpenStudy (anonymous):

is it right

OpenStudy (callisto):

I guess no?

OpenStudy (callisto):

Formula for integration by parts:\[\int u(x) v'(x) dx = u(x)v(x) - \int u'(x) v(x)dx\]

OpenStudy (anonymous):

same thing right

OpenStudy (callisto):

Now, you've got \[\int \frac{5}{2}xsin2xdx\]To simpler manipulations, pull out the constant \[=\frac{5}{2} \int xsin2xdx\]Now, what is u(x) and what is v'(x)?

OpenStudy (anonymous):

x

OpenStudy (callisto):

x is u(x) or v'(x)?

OpenStudy (anonymous):

u=x

OpenStudy (callisto):

*u(x) = x So, what is v'(x)?

OpenStudy (anonymous):

2cos 2x

OpenStudy (anonymous):

-

OpenStudy (callisto):

Nope..

OpenStudy (anonymous):

why

OpenStudy (callisto):